Consider a system of two identical particals. One of the particles is at rest and the other has an acceleration f.the centre of mass has acceleration
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
 
Let acm be the acceleration of the centre of mass of the two particles. Then,
                a_{cm}=\frac{m_{1}a_{1}+m_{2}a_{2}}{m_{1}+m_{2}}
Here since the particles are identical i.e, m1=m2=m(say)
          a_{cm}=\frac{m(a_{1}+a_{2})}{2m}=\frac{a_{1}+a_{2}}{2}               (1)
Taking a1=0(particle at rest) and a2= f, we get
        acm=f/2   
 
Regards
Arun (askIITians forum expert)